OpenStudy (anonymous):

How do you solve 2x/3 -1=3x-2/2?

OpenStudy (saifoo.khan):

\[\frac{2x}{3} -1 = \frac{3x-2}{2}\]Like that?

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

\[\frac{2x-3}{3} = \frac{3x-2}{2}\] Cross multiply, \[2(2x-3) = 3(3x-2)\] Can u solve further?

OpenStudy (anonymous):

i got x=0

OpenStudy (saifoo.khan):

Correct.

OpenStudy (anonymous):

thanks soo much!!

OpenStudy (saifoo.khan):

No problem!